OpenStudy (anonymous):

Graph the exponential function and state the domain and range. y = 5(x - 3)

OpenStudy (rina.r):

y=5x-15 y=15/5 y=3x i think

OpenStudy (tukitw):

Did you mean y = 5^(x - 3) instead? You did typed exponential...